Find the length of the third side of the right triangle. Give an exact answer and an approximation to three decimal places. I could not find a triangle to input into my question and I could not copy and paste from another program. The bottom is about on inch long with an 8 below the left side is straight up and down about 1/2 inch with a 6 beside with a small box in the corner. the right is the length of the bottom only coming down to meet the bottom with the letter c on top.
Answer by 24HoursTutor.com(40)

If your triangle looks like what I have imagined it to be from your description then this is the solution :
According to the Pythagoras theorem : a^2+ b^2 = c^2
Here a = 8, b = 6
Which means 8^2 + 6^2 = c^2
64 + 36 = c^2
100 = c^2
c = sqrt(100) = 10
Hence, the third side of the triangle is 10
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr